The density property insures that a rational number can be found between a pair of rational numbers. Find the specific rational number requested.
0.4 of the way between 15 and 50.
I get how to subtract 15 and 50 but i dont get where to go from (0.4(35)=14) i dont get how they got 14. Please help!

Amanda

Answer by stanbon(75887)

You need to go to the start point (15), then go 0.4 of (50-15) beyond it.
#= 15+0.4(50-15)
#=15+0.4(35)
#=15+14=29
Cheers,
Stan H.